OpenStudy (legomyego180):

Integrate:

OpenStudy (legomyego180):

\[\int\limits_{}^{}\frac{ 4x^2 }{ \sqrt{7+x^2} }\]

OpenStudy (legomyego180):

looks like trig sub, but the 7 under the radical makes me have doubts.

OpenStudy (legomyego180):

maybe IBP?

OpenStudy (thomas5267):

Use \(x=\sqrt{7}\tan(\theta)\)?

OpenStudy (legomyego180):

Ok, ill give it a shot

OpenStudy (legomyego180):

So... \[\int\limits_{}^{}\frac{ 4x^2 }{ \sqrt{7+x^2} }dx=\int\limits_{}^{}\frac{ 4*7\tan^2(\theta) }{ \sqrt{7+ (\sqrt{7} \tan(\theta))^2}} \sec^2(\theta)=\int\limits_{}^{}\frac{ 28\tan^2(\theta)\sec^2(\theta) }{ \sqrt{7(1+\tan^2(\theta))}}\] \[=\int\limits_{}^{}\frac{ 28\tan^2(\theta)\sec^2(\theta) }{ \sqrt{7(\sec^2(\theta)} }\]

OpenStudy (legomyego180):

So the point of trig sub is to get rid of the radical I think, but I cant see how it's helping me here yet

zepdrix (zepdrix):

Well if you split up the 7 and the secant,\[\large\rm =\frac{28}{\sqrt7}\int\limits \frac{\tan^2\theta \sec^2\theta~d \theta}{\sqrt{\sec^2\theta}}\]you should be able to take the root of that secant squared, ya?

zepdrix (zepdrix):

\[\large\rm =\frac{28}{\sqrt7}\int\limits\limits \frac{\tan^2\theta \sec^2\theta~d \theta}{\sec \theta}\]Cancel some stuff, and then sines and cosines, ya?

OpenStudy (legomyego180):

sines and cosines?

OpenStudy (legomyego180):

Im trying to do some weird u-sub

OpenStudy (legomyego180):

\[\frac{ 28 }{ \sqrt{7} }\int\limits_{}^{}\sec(\theta)(1-\sec^2(\theta)d(\theta)\]

zepdrix (zepdrix):

Ya.. looks like you'll have to do some secant reduction formula.. that's no fun :)

OpenStudy (legomyego180):

er, identities wrong. should be sec^2(theta)-1

zepdrix (zepdrix):

Distribute the secant,\[\large\rm =4\sqrt7\int\limits \sec^3\theta~d \theta-4\sqrt7\int\limits \sec \theta~d \theta\]

OpenStudy (legomyego180):

Sounds fishy, we never learned that. Are you sure I'm doing this problem the right way? Maybe by looking at the answer choices it will give us some insight?

zepdrix (zepdrix):

Recall that the integral of secant gives natural log of some business... So it seems like we're heading in the right direction. But yes, you can do it a different way if this feels uncomfortable. Integration by parts would certainly work. Never seen reduction formula for secant?\[\large\rm \int\limits \sec^{n}\theta~d \theta=\frac{1}{n-1}\sec^{n-2}\theta \tan \theta+\frac{n-2}{n-1}\int\limits \sec^{n-2}\theta~d \theta\]

OpenStudy (legomyego180):

nope, new to me

OpenStudy (legomyego180):

maybe calc III material?

zepdrix (zepdrix):

Naw it's calc 2 :D it's whatev though. Let's try Parts maybe? It might be a little confusing knowing how to break it up properly though.\[\large\rm \int\limits \frac{4x^2dx}{\sqrt{7+x^2}}\]Hmm well I know that if I integrate something like this, \(\large\rm \int \frac{x dx}{\sqrt{7+x^2}}\) then it should work out to something nice like \(\large\rm \sqrt{7+x^2}\) I think... So that's probably what we want for our dv, \(\large\rm dv=\frac{x~dx}{\sqrt{7+x^2}}\) and all the left over goes to the u, \(\large\rm u=4x\)

OpenStudy (legomyego180):

hm ok, I see what you did so far there

zepdrix (zepdrix):

\[\large\rm u=4x\qquad\qquad\qquad dv=\frac{x}{\sqrt{7+x^2}}dx\]Then,\[\large\rm du=4dx\qquad\qquad\qquad v=\sqrt{7+x^2}\]If you're unsure about that v, you can do a little side work, maybe a substitution to work it out.

OpenStudy (legomyego180):

\[4x \sqrt{7+x^2}-\int\limits_{}^{}4 \sqrt{7+x^2}\]

zepdrix (zepdrix):

Mmmm ok looks good so far

zepdrix (zepdrix):

And with a trig sub we'll end up at the same place as before. Ah that's no fun lol.

OpenStudy (legomyego180):

cant u sub it can you

zepdrix (zepdrix):

I'm not sure.. checking..

zepdrix (zepdrix):

No, doesn't seem so :\

OpenStudy (legomyego180):

Partial Fraction Decomposition doesnt seem to work either

zepdrix (zepdrix):

Why do all the options have 2x instead of 4x? Hmm

OpenStudy (legomyego180):

Im going to email my professor about it. Ill tell him what you said about secant reduction and see if there is an alternative method we learned Im not remembering, or maybe I skipped the day we discussed secant reduction lol

zepdrix (zepdrix):

It's a little confusing, so that might be a good idea :D Grr this 2x is really bugging me. I made a boo boo somewhere... I can't find it.

OpenStudy (legomyego180):

Ill let you know what he says! Thanks for the help anyway. I got some good practice on IBP and Trig sub from it regardless.

OpenStudy (legomyego180):

@zepdrix in case you were interested, my professor got back to me: I've included a picture of my work I sent to her. "First, you lost sqrt(7) next to sec T dT. Second, in your last line, you'll want to replace tan^2T with sec^2T-1. Then when you distribute that out, you'll have to use IBP to integrate the sec cubed part. It's a really ugly integral, look in the lectures on Powers of Trig. You can find what that integral equals."